Logbooks Lab Maintenance Evaporator_1 Evaporator_2 Laser cutter Target Production Test-Stand RH-ISAC RH-Cyclotron RH-Meson Hall RH-Beamlines RH-ARIEL
  Evaporator_2  Not logged in ELOG logo
Message ID: 599     Entry time: Tuesday, November 01, 2022, 17:35
Author: John Wong 
Target Material: TiC 
Target Oven W.O#:  
Source: TARGET BOAT 
Status: Done 
Subject: Np-TiC fibres systhesis 

Tuesday, November 01, 2022, 17:24

2 samples of np-TiO2/C fibres were produced using the e-spinner.  Each sample was wrapped with a piece of graphite sheet, and then loaded into a target container. 

Sample will be heated to 1700 C.

TGHT will be autoramped to 450 A; this should be close to 1600 C. (see elog#596)

Good vacuum: 3e-5 Torr

Bad vacuum: 5e-5 Torr

Omega slope: 0.99

Initiate autoramp: 1 A/min

Chamber has been pumped down below 4.2e-6 Torr.


 Tuesday, November 01, 2022, 22:46

 Pressure currently at 6.4e-7 Torr.

Initiated auto-ramping


 Wednesday, November 02, 2022, 09:37

TGHT was ramped up to 448A (readback 445-452A / 8.05V - 8.11V).  The voltage is much higher than the last process, see elog #596.

Temperature is much higher at this current, reading at 1870 - 1872 C. 

Will keep at this setting for now and continue to monitor.

See attachment#1 for the TGHT-Pressure plot.


 Wednesday, November 02, 2022, 21:35

Samples have been heated for over 16 hours; temp has been stabilized at 1875 C, vacuum at 7.3 e-7 Torr. 

Initiated cooling at 1A/min.

Chamber will be vented tomorrow morning.


 

Attachment 1: TiO2_C_fibres_carbonization_elog599.png  110 kB  Uploaded Wednesday, November 02, 2022, 10:44  | Show | Hide all
ELOG V2.9.2-2455